A solution is prepared by dissolving 6.00 g of an
I do not know even the formula for this problem
Answer #1

Osmotic pressure is a colligative property which will depend upon number of moles of solute as

osmotic pressure = concentration X R X T

Osmotic pressure = 0.75

concentration = Moles of solute / Volume of solution

= 0.0821 L atm / K mole

T = Temperature = 298 K

Concentration = P / RT = 0.75 / 0.0821 X 298 = 0.0306

0.0606 = Moles / 1 L

Moles = 0.0306

Mass / Mol wt = Mole = 0.0306 = 6g / Mol wt

Mol wt = 6 / 0.0306 = 196.07 g / mole